Demand: P=100-2Qd Supply: P=10+Qs What is the effect of a price floor at P=50

Demand: P=100-2Qd

Supply: P=10+Qs

What is the effect of a price floor at P=50
Answer #1

Market is at equilibrium when Qd = Qs

100 - 2Qd = 10 + Qs

90 = 3Q

Q = 30 units

P = 10 + 30 = $40

Now market price is $40.

A price floor of $50 is binding in this market as it lies above the market price

At this price, quantity demanded is (100 - 50)/2 = 25 units and quantity supplied is (50 - 10) = 40 units

Hence there is a surplus of unsold goods of 15 units when the price floor is applied at $50 in this market.